OpenStudy (anonymous):

7√3 - 4√6 + √48 - √54 What do I do since all the square roots are different?

OpenStudy (anonymous):

sqrt(48) = 4*sqrt(3) sqrt(54) = 3*sqrt(6)

OpenStudy (anonymous):

7√3 - 4√6 + √48 - √54 root(6) = root(3)*root(2) root(48) = root(3)*root(16) root(54) = root(3)*root(18)

OpenStudy (hoblos):

you try to find common sqrt √48 = 4√3 √54 = 3√6